OpenStudy (anonymous):

Evaluate g(x) = -x + 8 for g(-2)

OpenStudy (anonymous):

given g(x)=-x+8 so g(-2)=-(-2)+8 =2+8 =10

OpenStudy (anonymous):

Find the invese of f(x) = 4x + 12